# Structure of activity in multiregion recurrent neural networks

### v1

@@ Abstract
Neural circuits are composed of multiple regions, each with rich dynamics and engaging in communication with other regions. The combination of local, within-region dynamics and global, network-level dynamics is thought to provide computational flexibility. However, the nature of such multiregion dynamics and the underlying synaptic connectivity patterns remain poorly understood. Here, we study the dynamics of recurrent neural networks with multiple interconnected regions. Within each region, neurons have a combination of random and structured recurrent connections. Motivated by experimental evidence of communication subspaces between cortical areas, these networks have low-rank connectivity between regions, enabling selective routing of activity. These networks exhibit two interacting forms of dynamics: high-dimensional fluctuations within regions and low-dimensional signal transmission between regions. To characterize this interaction, we develop a dynamical mean-field theory to analyze such networks in the limit where each region contains infinitely many neurons, with cross-region currents as key order parameters. Regions can act as both generators and transmitters of activity, roles that we show are in conflict. Specifically, taming the complexity of activity within a region is necessary for it to route signals to and from other regions. Unlike previous models of routing in neural circuits, which suppressed the activities of neuronal groups to control signal flow, routing in our model is achieved by exciting different high-dimensional activity patterns through a combination of connectivity structure and nonlinear recurrent dynamics. This theory provides insight into the interpretation of both multiregion neural data and trained neural networks.
I Introduction
@@
A ubiquitous example of convergent evolution in nervous systems is the emergence of well-defined anatomical regions that interact with each other [1, 2, 3, 4]. Neurons show region-specific functional specialization to varying degrees [5, 6, 7, 8, 4], and this modularity is thought to underlie the flexible, adaptive capabilities of neural circuits [9, 10, 11]. Advances in neural-recording technologies have enabled the simultaneous monitoring of thousands of neurons across multiple brain regions in-vivo [12, 13, 14]. A general feature of these datasets is that while some signals are region-specific, others are broadly distributed across many regions. This structure raises a variety of questions about the mechanistic origin and functional consequences of the activity in multiregion neural circuits [15, 16, 17, 18].
@@
To answer such questions, a promising approach is to construct and analyze multiregion recurrent neural network models, trained either to perform cognitive tasks similar to those used in experiments [19, 20, 21, 22] or to generate recorded neural-activity data [8, 23]. These models have shed light on directed multiregion interactions involved in sensorimotor processing, context modulation, and changes in behavioral states [24].
Yet, in both neural circuits and the artificial counterparts constructed to understand them, the nature of multiregion interactions remains largely mysterious. For example, we lack an understanding of the synaptic connectivity that supports modular computations; even the basic principles of how within- and cross-region connectivity must differ to make a multiregion network distinct from a single large network are not well established. Functionally, how region-specific and network-wide dynamics coexist and interact, as well as the mechanisms by which networks flexibly route signals, are unclear.
@@
Here, we address these issues by analyzing a recurrent neural network model with multiple regions. Each region has a combination of random (unstructured) and low-rank (structured) connectivity, allowing isolated regions to generate both high-dimensional fluctuating activity and specific low-dimensional activity patterns [25, 26]. Inspired by experimental findings, the regions are connected by structured low-rank connectivity, enabling selective routing of activity patterns between regions while restricting the propagation of chaotic fluctuations.
In analogy with the structure of multiregion neural-activity data, this network has two forms of dynamics: high-dimensional fluctuations from disordered connectivity within regions, and low-dimensional signal transmission via low-rank cross-region connectivity. These forms of dynamics can be qualitatively different. For example, regions can show chaotic fluctuations while maintaining fixed points or limit cycles in their low-dimensional dynamics. Our analysis describes the coexistence and interaction of these two forms of dynamics.
In this model, regions function both as generators and transmitters of activity. When the intrinsic dynamics within a region grow strong or complex, the region may cease to transmit signals due to a conflict between signal generation and transmission. Our analysis characterizes this conflict and shows how taming within-region dynamics is crucial for communication at the network level. This behavior is reminiscent of the experimental finding that neuronal variability quenches at stimulus onset, suggesting that this quenching is a signature of of increased cross-region communication [27, 28].
@@
A key feature of our model is that signals are routed throughout the network not by activating or deactivating neurons in specific regions, but by shifting which subspaces of high-dimensional activity space, also called neural modes, are excited or unexcited through the combination of connectivity statistics and nonlinear recurrent dynamics [29]. The pattern of subspace excitation is determined by the geometric arrangement of the low-rank connectivity patterns and the strength of disordered connectivity. Thus, our model reflects a shift in perspective from earlier models of gating and routing in neural circuits, which emphasized single-neuron biophysical mechanisms [30], to a geometric, population-level view.
We analyze the network by deriving and solving a dynamical mean-field theory (DMFT). A central object in this DMFT is a third-order tensor that encodes the geometric arrangement of cross-region connectivity patterns. This effective-interaction tensor governs the evolution of order parameters that represent directed currents between pairs of regions, constituting the low-dimensional dynamics mentioned above. Additional order parameters (two-point functions) describe the high-dimensional dynamics. Thus, while neuron-to-neuron interactions are usually modeled as having a vector dynamics shaped by a matrix, in our framework, region-to-region interactions have a matrix dynamics shaped by a third-order tensor. The spectral properties of a matrix associated to this tensor often enable prediction of network behavior.
Our paper proceeds as follows. We start by introducing the connectivity model as an experimentally motivated way of balancing the extremes of fully specialized and homogeneous activity. We then derive DMFT equations for the network dynamics in the limit where each region has infinitely many neurons, for any intensive number of regions. To gain insight into the solutions of these DMFT equations, our analysis of them proceeds in stages. We start by considering symmetric effective interactions, leading to fixed-point solutions in the low-dimensional dynamics; then progress to include disorder; and finally consider general effective interactions, with the potential for low-dimensional limit-cycle solutions. The link between spectral structure of the effective-interaction tensor and the network behavior is particularly useful in this general case.
One challenge in analyzing this system is the tensorial nature of the effective interactions in the DMFT. We suggest that understanding the dynamics of large and complex neural circuits, characterized by significant modularity and hierarchy, will require grappling with multilinear interactions involving many indices. To this end, we have found the language of tensor diagrams from physics [31, 32] to be useful (Figs. 2,  10).
Overall, our study provides an analytically tractable model of multiregion interactions in high-dimensional neural systems based on combining an experimentally motivated form of connectivity with nonlinear recurrent dynamics. The DMFT we develop to describe this model, and the dynamic, subspace-based picture of routing that it provides, enhance our understanding of multiregion neural circuits. Furthermore, our work establishes a theoretical foundation for the implementation and interpretation of machine-learning tools that involve training multiregion recurrent networks to analyze neural data [9].
II Multiregion network model
@@
We study rate-based recurrent neural networks comprised of $R$ regions, with each region containing $N$ neuronal units. We keep $R$ finite and take $N\rightarrow\infty$. The preactivations of the neurons are denoted by $x_{i}^{\mu}(t)$, where ${\mu\in\{1,\ldots,R\}}$ specifies the region and ${i\in\{1,\ldots,N\}}$ specifies the within-region neuron. The post-activations are ${\phi_{i}^{\mu}(t)=\phi\left(x_{i}^{\mu}(t)\right)}$, where ${\phi(x)}$ is a pointwise nonlinearity. We use ${\phi(x)=\text{erf}(\sqrt{\pi}x/2)}$, which is similar to the more conventional $\tanh(x)$ but allows analytical evaluation of Gaussian integrals. The neurons interact through a synaptic coupling matrix $J^{\mu\nu}_{ij}$ according to
$(1+\partial_{t})x_{i}^{\mu}(t)=\sum_{\nu,j}J_{ij}^{\mu\nu}\phi_{j}^{\nu}(t).$ (1)
We assume that the couplings are randomly sampled from an ensemble, with zero mean for simplicity.
We first describe coupling ensembles that lead to either extreme functional specialization or homogeneity, both undesirable in a multiregion network. We then introduce a connectivity model that strikes a balance between these extremes through selective signal routing.
Consider first a network with no cross-region couplings and independently and identically distributed (i.i.d.) within-region couplings,
$J^{\mu\nu}_{ij}=\delta^{\mu\nu}\chi^{\mu}_{ij},$ (2)
where $\chi^{\mu}_{ij}$ has zero mean and $\left\langle(\chi^{\mu}_{ij})^{2}\right\rangle_{\bm{J}}=(g^{\mu})^{2}/N$ (throughout, $\left\langle\cdot\right\rangle_{\bm{J}}$ will denote an average with respect to the coupling ensemble in question). This represents $R$ independent versions of the classic network model of Sompolinsky et al. [33], potentially with different $g^{\mu}$ values across $\mu$. In this case, each region $\mu$ transitions from quiescence to chaos at a critical coupling variance of $g^{\mu}=1$. This results in a trivial form of regional specialization where the regions do not communicate.
For nontrivial interregion dynamics, the cross-region terms in $J^{\mu\nu}_{ij}$ must be nonzero and have statistics that depend on $(\mu,\nu)$–in particular, the statistics for $\mu=\nu$ and $\mu\neq\nu$ should be different. Along these lines, Aljadeff et al. [34] considered independently but not identically distributed couplings of the form
$J^{\mu\nu}_{ij}=\chi^{\mu\nu}_{ij},$ (3)
where $\chi^{\mu\nu}_{ij}$ has zero mean and $\left\langle(\chi^{\mu\nu}_{ij})^{2}\right\rangle_{\bm{J}}=G^{\mu\nu}/RN$, with $G^{\mu\nu}$ being a matrix of variance parameters. A transition to chaos occurs when $G^{\mu\nu}$ acquires an eigenvalue with real part larger than unity. However, despite having region-dependent statistics, such couplings do not produce regionally specialized activity. Instead, chaotic activity is distributed globally across regions. For example, within- and cross-region pairwise correlations between neuronal activities are not differentiated in their scaling, which is ${\sim 1/\sqrt{N}}$ in both cases. Additionally, region-specific single-neuron two-point functions differ only by linear transformations. We will later show that the global propagation of chaotic fluctuations in this network is due to its high-dimensional cross-region connectivity.
Figure 1: Top: Schematic of the synaptic connectivity model. Different regions, each of which has “random plus rank-one” connectivity, are linked via rank-one matrices representing communication subspaces. In this network of $R=4$ regions, we highlight the rank-one and disordered couplings in region $\mu$, as well as the structured couplings to and from region $\nu$. Rank-one connections are defined through the outer product of vectors $n_{i}^{\mu\nu}$ and $m_{i}^{\nu\rho}$. Bottom: Effective-interaction tensor $T^{\mu\nu\rho}$, which encodes the geometric arrangement of the connectivity patterns and determines the dynamics of region-to-region currents in the mean-field picture. The entry $T^{\mu\nu\rho}$ measures the overlap between $n_{i}^{\mu\nu}$ and $m_{i}^{\nu\rho}$, contracted over the neuron index $i$.
Our approach to generate network behavior between these extremes is rooted in experiments demonstrating that cortical areas interact through so-called communication subspaces. In particular, Semedo et al. [35] found that only a low-dimensional subspace of neural activity in primary visual cortex, V1, shows correlation with activity in V2, and this subspace is distinct from the subspace that captures the most variance of the activity in V1. This can be explained by the synaptic connectivity matrix from V1 to V2 having low-rank structure. In this case, whether activity in V1 is transmitted to V2 depends on the degree of alignment of the activity with the column space of this matrix. That is, an activity-dependent bottleneck limits communication between regions. This study was followed by others demonstrating communication subspaces relevant to visual processing [36], motor control [37, 38], attention [39], audition [22], and brain-wide activity during unstructured behavior [40].
We operationalize this idea by considering couplings of the form
$J^{\mu\nu}_{ij}=\delta^{\mu\nu}\chi^{\mu}_{ij}+\frac{1}{N}m^{\mu\nu}_{i}n^{\mu\nu}_{j},$ (4)
where $\left\langle(\chi^{\mu}_{ij})^{2})\right\rangle_{\bm{J}}=(g^{\mu})^{2}/N$. Here, the synaptic connections within region $\mu$ are the sum of random disordered couplings, $\chi^{\mu}_{ij}$, and a rank-one matrix, $m_{i}^{\mu\mu}n_{j}^{\mu\mu}/N$. These terms represent the combination of disorderly and structured recurrent circuitry within brain regions [25]. Meanwhile, to model communication subspaces, the cross-region couplings from region $\nu$ to region $\mu$, for $\mu\neq\nu$, are a rank-one matrix $m_{i}^{\mu\nu}n_{j}^{\mu\nu}/N$ (Fig. 1, top). Note that we study the case where there is no additional disorder in the cross-region connectivity, but we provide DMFT equations incorporating such disorder in Sec.VII (see Eq. 38). The “degree of alignment of the activity with the column space” of a cross-region connectivity matrix, mentioned in the previous paragraph, is given by the inner product of an activity vector $\phi_{i}^{\nu}$ and a connectivity pattern $n^{\mu\nu}_{i}$; these inner products are order parameters in the DMFT.
@@
In this model, there are $2R^{2}$ vector components for each neuron index $i$, namely, ${\{m_{i}^{\mu\nu}\}_{\mu,\nu=1}^{R}\cup\{n_{i}^{\mu\nu}\}_{\mu,\nu=1}^{R}}$. These components are sampled from a zero-mean multivariate Gaussian, i.i.d. across $i$. The second-order statistics of this Gaussian are defined by the tensors
@@
Specification of the remaining second-order statistics is not required for the mean-field description of the model. However, to sample the vectors defining the low-rank part of the couplings, we must specify $\left\langle n_{i}^{\mu\nu}n_{i}^{\rho\nu}\right\rangle_{\bm{J}}$. We set this proportional to $\delta^{\mu\rho}$ with a positive scale factor large enough so that the full covariance matrix is positive-definite. This scale-factor degree of freedom allows any pair of tensors $T^{\mu\nu\rho}$ and $U^{\mu\nu\rho}$, the latter being symmetric and positive-semidefinite in $(\nu,\rho)$ for all $\mu$, to define a valid distribution over couplings.
As $N\rightarrow\infty$, these tensors are equivalently given by inner products between vectors,
$T^{\mu\nu\rho}$ $=\frac{1}{N}\sum_{i}{n_{i}^{\mu\nu}m_{i}^{\nu\rho}},$ (6a)
$U^{\mu\nu\rho}$ $=\frac{1}{N}\sum_{i}{m_{i}^{\mu\nu}m_{i}^{\mu\rho}}.$ (6b)
Hence, $T^{\mu\nu\rho}$ and $U^{\mu\nu\rho}$ represent the geometric arrangement of structured low-rank connectivity patterns (Fig. 1, bottom)  [^1] Specifying the covariances $\left\langle n_{i}^{\cdot\mu}m_{i}^{\nu\cdot}\right\rangle_{\bm{J}}$, $\left\langle m_{i}^{\mu\cdot}m_{i}^{\nu\cdot}\right\rangle_{\bm{J}}$, and $\left\langle n_{i}^{\cdot\mu}n_{i}^{\cdot\nu}\right\rangle_{\bm{J}}$ for $\mu\neq\nu$ is unnecessary for both the mean-field description and for sampling couplings. This is because such covariances, when interpreted as vector contractions, are not physically valid; the relevant vectors only have the same number of components $N$ in our model for simplicity, but in general their numbers of components could differ across regions by intensive ratios..

### v3

@@ Abstract
Neural circuits comprise multiple interconnected regions, each with complex dynamics. The interplay between local and global activity is thought to underlie computational flexibility, yet the structure of multiregion neural activity and its origins in synaptic connectivity remain poorly understood. We investigate recurrent neural networks with multiple regions, each containing neurons with random and structured connections. Inspired by experimental evidence of communication subspaces, we use low-rank connectivity between regions to enable selective activity routing. These networks exhibit high-dimensional fluctuations within regions and low-dimensional signal transmission between them. Using dynamical mean-field theory, with cross-region currents as order parameters, we show that regions act as both generators and transmitters of activity—roles that are often in tension. Taming within-region activity can be crucial for effective signal routing. Unlike previous models that suppressed neural activity to control signal flow, our model achieves routing by exciting different high-dimensional activity patterns through connectivity structure and nonlinear dynamics. Our analysis offers insights into multiregion neural data and trained neural networks.
1 Introduction
@@
A striking example of convergent evolution in nervous systems is the emergence of well-defined anatomical regions that interact with one another [1, 2, 3, 4]. Recent advances in neural-recording technologies have enabled simultaneous monitoring of thousands of neurons across multiple brain areas in vivo [5, 6, 7, 8]. These studies reveal that neurons exhibit varying degrees of regional specialization in their activities [9, 10, 11, 4]. This regional specialization, balanced with cross-region interactions, is believed to underlie the flexible, adaptive capabilities of neural circuits [12, 13, 14]. Modern neural datasets thus reveal an intricate interplay between region-specific and broadly distributed signals.
@@
These datasets raise fundamental questions about the origins and functions of multiregion neural activity [15, 16, 17, 18]. To address them, researchers have trained multiregion recurrent neural network models, either to perform cognitive tasks [19, 20, 21, 22] or to generate recorded neural data [23, 24]. These models have shed light on directed multiregion interactions involved in sensorimotor processing, context modulation, and changes in behavioral states [25].
However, in both real neural circuits and their artificial counterparts, the nature of multiregion interactions remains largely mysterious. In particular, we lack understanding of the connectivity supporting modular computations and the mechanisms of flexible signal routing. The coexistence and interaction of region-specific and network-wide dynamics are also unclear.
@@
To address these challenges, we analyze a recurrent network model with multiple regions. Each region has a combination of random and low-rank connectivity, generating both high-dimensional fluctuations and specific low-dimensional patterns [26, 27]. We connect regions using low-rank connectivity, enabling selective routing of low-dimensional signals between regions.
Due to its nonlinear dynamics and multiregion connectivity structure, this model produces an extremely rich and broad array of dynamic states depending on the connectivity. We develop an analytical theory of this multiregion activity structure by deriving and solving dynamical mean-field theory (DMFT) equations for the network in the limit where each region has infinitely many neurons for any finite number of regions. Given the complexity of the resulting DMFT equations, we solve them in stages of increasing complexity: first considering symmetric effective interactions leading to fixed-point solutions in the low-dimensional dynamics, then progressing to include disorder. Finally, we examine general effective interactions with the potential for limit-cycle solutions, requiring numerical solution.
Our analysis reveals two key ideas, each supported by various specific results:
Key idea 1: Regions serve dual roles as generators and transmitters of activity, with an inherent tension between these functions. When the intrinsic dynamics within a region become too strong or complex, the region’s ability to transmit signals is compromised. Our analysis characterizes this conflict and demonstrates how taming within-region dynamics is crucial for network-level communication.
@@
Key idea 2: Signal routing throughout the network is achieved by shifting which subspaces of high-dimensional activity space are excited or unexcited through the interplay of connectivity statistics and nonlinear recurrent dynamics. The subset of subspaces that are excited depends on the geometric arrangement of low-rank connectivity patterns and the strength of disordered connectivity. Our approach complements earlier models of gating and routing in neural circuits, which emphasized single-neuron biophysical mechanisms such as neuromodulation, inhibition, or gain modulation [28], by developing a geometric, population-level view of information flow.
Overall, our work provides a theoretical framework for understanding the interplay between regional specialization and multiregion interactions in neural circuits, offering insights into the mechanisms underlying flexible signal routing and modular computations in the brain.
2 Multiregion Network Model
Here, we present the multiregion network model, first describing its dynamics and then its connectivity.
2.1 Dynamics
@@
We study rate-based (non-spiking) recurrent neural networks comprising $R$ regions, each containing $N$ neurons. We consider a finite number of regions $R$ and take the limit $N\rightarrow\infty$, corresponding to a small or moderate number of regions, each with a large number of neurons. The preactivations of the neurons, analogous to membrane potentials, are denoted by $x_{i}^{\mu}(t)$, where $\mu\in\{1,\ldots,R\}$ specifies the region and $i\in\{1,\ldots,N\}$ specifies the within-region neuron. The activations, analogous to firing rates, are given by $\phi_{i}^{\mu}(t)=\phi(x_{i}^{\mu}(t))$, where $\phi(x)=\text{erf}(\sqrt{\pi}x/2)$ is a pointwise nonlinearity that is linear for small $|x|$ and saturates at ±$\pm 1$ for large $|x|$. Neurons interact through a synaptic coupling matrix $J^{\mu\nu}_{ij}$ according to:
$\frac{dx_{i}^{\mu}(t)}{dt}=-x_{i}^{\mu}(t)+\sum_{\nu=1}^{R}\sum_{j=1}^{N}J_{ij}^{\mu\nu}\phi_{j}^{\nu}(t).$ (1)
2.2 Connectivity
The connections within each region $\mu$ are dense and consist of the sum of random disordered couplings, $\chi^{\mu}_{ij}$, and a rank-one matrix, as investigated by Mastrogiuseppe and Ostojic [26]. This rank-one matrix is defined as the outer product of vectors $\bm{m}^{\mu\mu}$ and $\bm{n}^{\mu\mu}$. Connections between pairs of regions, such as from region $\nu$ to $\mu$, are represented by additional rank-one matrices formed by outer products of vectors $\bm{m}^{\mu\nu}$ and $\bm{n}^{\mu\nu}$ (Fig. 1a). The synaptic coupling matrix is thus expressed as:
$J^{\mu\nu}_{ij}=\delta^{\mu\nu}\chi^{\mu}_{ij}+\frac{1}{N}m^{\mu\nu}_{i}n^{\mu\nu}_{j}.$ (2)
Each element of $\chi^{\mu}_{ij}$ is sampled independently from a zero-mean Gaussian with variance $(g^{\mu})^{2}/N$. This $1/\sqrt{N}$ scaling of the disordered couplings ensures that the eigenspectrum of $\chi^{\mu}_{ij}$ remains independent of network size for large $N$.
@@ 2.2 Connectivity
For tractability, we assume that the components of the vectors $\bm{m}^{\mu\nu}$ and $\bm{n}^{\mu\nu}$ are zero-mean random variables drawn from a multivariate Gaussian. Specifically, for each neuron in the network, such as for neuron $i$ in region $\nu$, there are $2R$ jointly sampled components: $\{n_{i}^{\mu\nu}\}_{\mu=1}^{R}\cup\{m_{i}^{\nu\rho}\}_{\rho=1}^{R}$. To define the second-order statistics of these components, we introduce the tensors:
@@
Our analysis will demonstrate that specifying the remaining second-order statistics, $\left\langle n_{i}^{\mu\nu}n_{i}^{\rho\nu}\right\rangle_{\bm{J}}$, is not necessary to study the dynamics in the limit $N\to\infty$. However, to sample the vectors defining the low-rank part of the couplings, we must specify $\left\langle n_{i}^{\mu\nu}n_{i}^{\rho\nu}\right\rangle_{\bm{J}}$. We set this proportional to $\delta^{\mu\rho}$ with a scale factor large enough to ensure that the overall covariance matrix of vector components is positive-definite. As $N\to\infty$, these tensors can equivalently be expressed by the normalized overlaps or inner products:
$T^{\mu\nu\rho}$ $=\frac{1}{N}\sum_{i=1}^{N}{n_{i}^{\mu\nu}m_{i}^{\nu\rho}},$ (4a)
$U^{\mu\nu\rho}$ $=\frac{1}{N}\sum_{i=1}^{N}{m_{i}^{\mu\nu}m_{i}^{\mu\rho}}.$ (4b)
Thus, $T^{\mu\nu\rho}$ and $U^{\mu\nu\rho}$ encode the geometric arrangement of connectivity patterns (Fig. 1a, bottom), providing a concise representation of the network’s structure. When showing simulation results, we will consider only large networks where the particular realization of connectivity is not significant, and the system behavior is controlled by $g^{\mu}$, $T^{\mu\nu\rho}$, and $U^{\mu\nu\rho}$.
Table 1 summarizes the variables and notation used throughout this article.
Figure 1: (a) Top: Schematic of the synaptic connectivity model. Different regions, each with “random plus rank-one” connectivity, are linked via rank-one matrices representing communication subspaces. In this network of $R=4$ regions, we highlight the rank-one and disordered couplings in region $\mu$, as well as the structured couplings to and from region $\nu$. Rank-one connections are defined through the outer product of vectors $\bm{m}^{\mu\nu}$ and $\bm{n}^{\mu\nu}$. Bottom: Tensor $T^{\mu\nu\rho}$, which encodes the geometric arrangement of the connectivity patterns and determines the dynamics of region-to-region currents in the mean-field picture. (b) Anatomical bottleneck or effective bottleneck implementing a rank-one connectivity matrix between regions $\nu$ and $\mu$. The dashed circle represents a linear neuron with fast timescale.
3 Biological Motivations and Assumptions
In constructing this model, we aimed to incorporate sufficient biological detail to capture nontrivial phenomena while maintaining analytical tractability. In this section, we elucidate the biological foundations of our model, outlining its underlying assumptions and limitations, first addressing the dynamics and then the connectivity.
3.1 Dynamics: Motivation and Assumptions
The complexity in our network model’s dynamics, compared to linear networks that can simply be diagonalized, stems from the nonlinear activations of individual neurons. This nonlinearity is inspired by the transformation of input currents into spike trains by real neurons. While our model captures this crucial aspect, it does not account for other features of cortical circuits, such as distinct excitatory and inhibitory populations (i.e., Dale’s law), sparse connectivity, and nonnegative firing rates.
This level of abstraction mirrors that used in the seminal work of Sompolinsky et al. [29], which described chaotic activity arising from strong random connectivity. Indeed, our multiregion model reduces to $R$ independent samples of this model when the structured low-rank couplings are set to zero. In this special case, each disconnected region transitions from quiescence to high-dimensional chaos at a critical coupling variance, defined by $g^{\mu}=1$.
Our use of this level of abstraction is supported by recent studies demonstrating that network models incorporating the biological features we omitted (i.e., nonnegative rates or spikes, sparse connections, and excitatory-inhibitory populations) can exhibit equivalent dynamical regimes. This equivalence has been observed both for disordered couplings, where the same transition to chaos occurs [30, 31], and for low-rank couplings [32, 33].
3.2 Connectivity: Motivation and Assumptions
We use rank-one matrices to model structured connectivity both within and between regions, based on separate experimental observations for each type of connectivity.
Within-region recordings show that neural activity during tasks often lies on a low-dimensional manifold [34, 26]. Rank-one connectivity can generate arbitrary one-dimensional dynamics [35], serving as a starting point for modeling structured low-dimensional activity. Many standard neural-network models, including Hopfield networks [36], ring attractors [37], and autoencoders [38], use low-rank connectivity. Furthermore, our model combines rank-one and disordered within-region connectivity. As shown by Mastrogiuseppe and Ostojic [26], such networks can produce chaotic activity, fixed points, or both, depending on the relative strengths of rank-one vs. disordered connectivity.
Cross-region rank-one connections are based on observed communication subspaces between cortical areas. In particular, Semedo et al. [39] found that only a low-dimensional subspace of V1 activity, distinct from the subspace capturing most V1 variance, correlates with activity in V2. Similar communication-subspace structure has been identified in visual processing [40], motor control [41, 42], attention [43], audition [22], and brain-wide activity [44]. Low-rank cross-region connectivity offers a simple explanation for these subspaces, but of course is not the only explanation. Alternative hypotheses, such as global fluctuations or shared input, were considered less likely based on anatomy, spatial selectivity, and persistence under anesthesia by the authors of the original study (in visual cortex). Here, we adopt low-rank connectivity for its simplicity, data compatibility, and, as we discuss in the next section, functional utility.
Biologically, low-rank cross-region connectivity, which acts as a type of bottleneck, can be implemented either anatomically or effectively (Fig. 1b; [45, 26]). An anatomical bottleneck would involve a set of intermediary neurons between two areas (Fig. 1b, top). These neurons, assumed to be linear with fast time constants, would read out activity from the source region and broadcast it to the target region [46]. This framework also accommodates thalamocortical loops as anatomical bottlenecks between cortical regions (this complements existing models where thalamic nuclei create loops within a cortical area; such loops can be selectively modulated via basal-ganglia inhibition, controlling inter-region communication). Alternatively, an effective bottleneck would arise from direct, monosynaptic connections between source and target regions with a low-rank structure (Fig. 1b, bottom). A simple example of this occurs when all connections from a source to a target region have the same strength and sign, corresponding to a rank-one matrix that is sensitive only to the mean activity of the source region.
Under the interpretation of an effective bottleneck, the rank-one constraint results in a synaptic coupling from a neuron in region $\nu$ to a neuron in region $\mu$ that is proportional to the product of two scalar variables: $n^{\mu\nu}_{i}$ and $m^{\mu\nu}_{j}$. These variables are associated with the emitter and receiver populations, respectively. Such couplings, expressed as products of pre- and postsynaptic terms, arise naturally in neuroscience as a consequence of Hebbian plasticity.
Finally, while we use rank-one matrices, a more realistic model might involve higher-rank matrices, or matrices with smoothly decaying singular values. We find that even rank-one matrices induce rich multiregion activity structure, providing an adequate starting point.
3.3 Functional Significance of Low-Rank Cross-Region Connectivity
A rank-one connectivity matrix implements an activity-dependent bottleneck: the transmission of activity from source region $\nu$ to target region $\mu$ depends on the alignment of activity in $\nu$ with the row space of the connecting low-rank matrix. This row space, given by the span of $\bm{n}^{\mu\nu}$, represents the communication subspace in our model. The bottleneck then projects this filtered activity into target region $\mu$ through the column space of the matrix, given by the span of $\bm{m}^{\mu\nu}$.